OpenStudy (anonymous):

Suppose P9,500 is invested at 6% interest compounded annually. Approximately how long will it take for the amount in the account to reach P12,000?

OpenStudy (anonymous):

P12,000 = P9,500 *1.6X

OpenStudy (phi):

use \[ P= P_0 (1.06)^{t} \] where P is 12000, P0 is 9500 and t is the number of years: \[ 12000 = 9500 (1.06)^t \] to solve , divide both sides by 9500 \[ \frac{12000}{9500} = \frac{9500}{9500} (1.06)^t \] or \[ 1.06^t = 1.26316 \] take the log of both sides: \[ \log(1.06^t) = \log(1.26316) \] use the rule \[ \log(a^b) = b\ \log(a) \] can you finish ?